OpenStudy (anonymous):

ok, the series converges. \[\sum_{n=1}^{infty} (-1)^{n}/ n 5^{n}\] How many term of the series do we need to add in order to find the sum to the indicated accuracy. (|error| <0.0001)

OpenStudy (anonymous):

0.22266666666666666666666666 is the addition of 3 terms

OpenStudy (anonymous):

how do i find the error again? I would divide the above number by....?

OpenStudy (experimentx):

is it \[ \sum_{n=1}^\infty \frac{(-1)^n}{n \times 5^n}\]

OpenStudy (experimentx):

If it's either of them individually, then they have standard values.

OpenStudy (experimentx):

for (|error| <0.0001) ... evaluating upto 4 terms would be sufficient.

OpenStudy (anonymous):

ohh..I didn't see the replies....THANKS!!!

OpenStudy (anonymous):

In an alternating series the error is less than the first unused term. So @experimentX is right since \[ \frac{1}{5. 5^5}=0.000064 < .0001 \]

OpenStudy (anonymous):

thanks!

OpenStudy (experimentx):

yw

OpenStudy (anonymous):